# Unifying KV Cache Compression for Large Language Models with LeanKV

### v2

@@ 7.3 System Performance
In this section, we provide a comprehensive evaluation of LeanKV’s performance, comparing it against state-of-the-art systems including vLLM, Quest, and SnapKV across various models. All experiments are conducted on a machine equipped with eight NVIDIA L40 GPUs. For models that fit within the memory of a single GPU, such as Llama3-8B and Qwen2.5-7B, we use one GPU. For larger models, we parallelize the computation across multiple GPUs: four GPUs for Llama3-70B, and two GPUs for Qwen2.5-32B and QwQ-32B. Our evaluation focuses on latency breakdown, attention kernel speedup, and end-to-end throughput to demonstrate the efficiency and scalability of LeanKV.
Memory Management Overhead. Figure 11 compares LeanKV’s on-GPU parallel KV compaction with an alternative implementation that performs multi-threaded memory management on the CPU. Evaluated with a sequence length of 1024 tokens across different batch sizes, LeanKV reduces memory management latency by up to three orders of magnitude relative to the on-CPU approach. For one entire inference step, LeanKV significantly outperforms the on-CPU approach, particularly in the generation phase, where the memory management overhead on the CPU far exceeds the model execution time. Therefore, LeanKV’s on-GPU parallel memory management is essential for ensuring that the performance benefits of KV cache compression are fully realized without being overshadowed by memory management overhead.
Figure 12 further illustrates the latency breakdown of LeanKV during one inference step. The memory management overhead is remarkably low thanks to LeanKV’s on-GPU parallel KV compaction, contributing less than 0.2% of the total latency in the prompt phase and under 0.9% in the generation phase. Model execution dominates the latency in both phases, accounting for 96–97% of the latency in the prompt phase and 92–93% in the generation phase. Notably, as the batch size increases, the percentage of time spent on model execution rises slightly, reflecting the scalability of LeanKV.
@@
Figure 12: Latency breakdown of LeanKV.
Attention Kernel Speedup. Figure 13 shows the speedup of the LeanKV’s custom attention kernel against vLLM under different quantization configurations. LeanKV achieves a near-linear speedup proportional to the reduction in KV cache size. For example, with K8V8, which halves the KV cache size relative to FP16 datatype, the theoretical speedup is 2$\times$, and LeanKV achieves 1.7$\times$. The slight gap is primarily due to the overhead of accessing quantization metadata and performing dequantization. Additionally, LeanKV achieves greater speedups on longer sequences, indicating that its bandwidth optimization techniques become increasingly effective as the sequence length grows.
Figure 13: Attention kernel speedup.
End-to-End Throughput. Figure 14 presents the end-to-end throughput and achieved batch sizes of LeanKV compared to vLLM, Quest, and SnapKV. The maximum generation length is set to 16K tokens for QwQ-32B, 8K for Qwen2.5-32B, and 4K for the other models. We evaluate on 1000 sequences sampled from the MATH dataset [34], which naturally elicits chain-of-thought reasoning and typically leads to long generations reaching the specified limit. Compression thresholds for LeanKV are adopted from Figure 9. Across all models, LeanKV consistently achieves higher throughput than prior systems. Notably, LeanKV achieves a remarkable 5.4×5.4$\times$5.4 × higher throughput over vLLM on the QwQ-32B thinking model, while Quest and SnapKV achieve only 1.6×1.6$\times$1.6 × and 1.8×1.8$\times$1.8 × speedups, respectively. This result highlights the superior efficiency of LeanKV in processing long sequences, making it particularly well-suited for tasks that require extended reasoning.
The throughput improvements are directly correlated with the larger batch sizes supported by LeanKV, enabled by its KV cache compression techniques. For instance, with the QwQ-32B model, LeanKV sustains a batch size of 15.9, far exceeding vLLM’s batch size of 2.7. The increased batch size enables LeanKV to maximize the utilization of GPU compute capacity, leading to higher throughput. In contrast, Quest supports the same batch size as vLLM because it retains the entire KV cache without reducing memory usage. Its speedup comes primarily from faster attention computation, as it processes only a subset of tokens deemed important. However, the process of estimating token importance incurs additional overhead, limiting the overall efficiency of Quest.
@@ 8 Conclusion
We propose LeanKV, a framework to enhance LLM serving efficiency by exploiting three levels of differentiation in KV cache: differentiated precision for keys and values, hierarchical compression based on varying token importance, and per-head dynamic sparsity. At the core of LeanKV is the parallel KV compaction technique that efficiently handles irregular memory requirements across requests and attention heads, effectively translating memory savings into performance gains. Our evaluation shows that LeanKV is able to compress the KV cache by 2.7×2.7$\times$2.7 × to 5.7×5.7$\times$5.7 × without compromising accuracy on complex workloads that require sophisticated reasoning and long-generation capabilities, and enhances throughput by 1.9×1.9$\times$1.9 × to 5.4×5.4$\times$5.4 ×.

### v3

@@ 7.3 System Performance
In this section, we provide a comprehensive evaluation of DiffKV’s performance, comparing it against state-of-the-art systems including vLLM, Quest, and SnapKV across various models. For models that fit within the memory of a single GPU, such as Llama3-8B and Qwen2.5-7B, we use one GPU. For larger models, we parallelize the computation across multiple GPUs: four GPUs for Llama3-70B, and two GPUs for Qwen2.5-32B and QwQ-32B. Our evaluation focuses on latency breakdown, attention kernel speedup, and end-to-end throughput to demonstrate the efficiency and scalability of DiffKV.
Memory Management Overhead. Figure 13 compares Diff-KV’s on-GPU parallel KV compaction with an alternative implementation that performs multi-threaded memory management on the CPU. Evaluated with a sequence length of 1024 tokens across different batch sizes, DiffKV reduces memory management latency by up to three orders of magnitude relative to the on-CPU approach. For one entire inference step, DiffKV significantly outperforms the on-CPU approach, particularly in the generation phase, where the memory management overhead on the CPU far exceeds the model execution time. Therefore, DiffKV’s on-GPU parallel memory management is essential for ensuring that the performance benefits of KV cache compression are fully realized without being overshadowed by memory management overhead.
Figure 14 further illustrates the latency breakdown of Diff-KV during one inference step. The memory management overhead is remarkably low thanks to DiffKV’s on-GPU parallel KV compaction, contributing less than 0.2% of the total latency in the prompt phase and under 0.9% in the generation phase. Model execution dominates the latency in both phases, accounting for 96–97% of the latency in the prompt phase and 92–93% in the generation phase. Notably, as the batch size increases, the percentage of time spent on model execution rises slightly, reflecting the scalability of DiffKV.
@@
Figure 14: Latency breakdown of DiffKV.
[…]
(a) Attention kernel
(b) One batch end-to-end
Figure 15: Latency speedup of DiffKV against vLLM.
Throughput Speedup. Figure 17 presents the end-to-end throughput and achieved batch sizes of DiffKV compared to vLLM, pruning-based approaches including Quest and SnapKV, as well as quantization-based methods including Atom (4-bit) and KIVI (2-bit). Note that these pruning- and quantization-based baselines incur larger accuracy degradation than DiffKV, as shown in Table 1. The maximum generation length is set to 16K tokens for QwQ-32B, 8K for Qwen2.5-32B, and 4K for the other models. We evaluate on 1000 sequences sampled from the MATH dataset [41], which elicits chain-of-thought reasoning and typically leads to long generations reaching the specified limit. Compression thresholds for DiffKV are adopted from Figure 10. Across all models, DiffKV consistently achieves higher throughput than prior systems. Notably, DiffKV achieves a remarkable $5.4\times$ higher throughput over vLLM on the QwQ-32B thinking model, while Quest, SnapKV, Atom, and KIVI achieve $1.6\times$, $1.8\times$, $2.1\times$, and $3.4\times$ speedups, respectively. This result highlights the superior efficiency of DiffKV in processing long sequences, making it particularly well-suited for tasks that require extended reasoning.
The throughput improvements are directly correlated with the larger batch sizes supported by DiffKV, enabled by its KV cache compression techniques. For instance, with the QwQ-32B model, DiffKV sustains a batch size of 15.9, far exceeding vLLM’s batch size of 2.7. The increased batch size enables DiffKV to maximize the utilization of GPU compute capacity, leading to higher throughput. In contrast, Quest supports the same batch size as vLLM because it retains the entire KV cache without reducing memory usage. Its speedup comes primarily from faster attention computation, as it processes only a subset of tokens deemed important. However, the process of estimating token importance incurs additional overhead, limiting the overall efficiency of Quest. Note that although Atom and KIVI substantially increase the batch size, approaching or even exceeding DiffKV, their throughput gains are not proportional. This is because both methods are built on the HuggingFace Transformers library, which incurs higher framework overhead than vLLM and DiffKV, and lacks high-performance GPU kernels that fuse dequantization with attention computation, as implemented in DiffKV. These results underscore that custom GPU kernels and an efficient runtime are essential for effectively translating KV cache compression into throughput gains.
Serving Dynamic Workloads. We evaluate DiffKV under dynamic workloads and compare it to vLLM [39]. Following the methodology of vLLM, we inject requests according to a Poisson arrival process and sweep the request rate. Figure 16 reports the average per-token latency (including both queuing and LLM processing time) for Llama3-8B and Qwen2.5-32B across a range of request rates. DiffKV consistently achieves lower latency and sustains higher loads before queuing delays grow sharply. These improvements stem from DiffKV’s reduced KV cache footprint, which enables larger batch sizes, and its parallel on-GPU KV compaction, which keeps memory management overhead low even under bursty traffic.
Figure 16: Comparison of average latency between DiffKV and vLLM under dynamic workloads.
@@ 8 Conclusion
We proposed DiffKV, which enhances LLM serving efficiency by exploiting three levels of differentiation in KV cache: differentiated precision for keys and values, hierarchical compression based on token importance, and per-head dynamic sparsity. At the core of DiffKV is the parallel KV compaction technique that efficiently handles irregular memory requirements across requests and attention heads, effectively translating memory savings into performance gains. Our evaluation demonstrated that DiffKV is able to compress the KV cache by $2.7\times$ to $5.7\times$ with near-lossless accuracy, even on thinking models and complex workloads that require sophisticated reasoning and long-generation capabilities, and enhances throughput by $1.9\times$ to $5.4\times$, outperforming prior KV cache compression methods.
9 Acknowledgment
We sincerely thank our shepherd, Tim Harris, for his thoughtful feedback and careful guidance throughout the paper shepherding process. We are also thankful to Mingzhe Hao, Shuang Chen, other friends, and the anonymous reviewers for their insightful comments on earlier versions of this manuscript.